OpenStudy (anonymous):

CAN SOMEONE HELP ME?!?!?!?! Question 4. 4. Write the product as a trinomial. (3r − 4)(r + 8) 3r2+ 20r − 32 3r2 − 28r − 32 3r2 − 32 4r2 + 4r − 32

OpenStudy (anonymous):

3r2 + 24r - 4r - 32 = 3r2 + 20r -32

OpenStudy (anonymous):

How did you get that?! Cause I got 3r2+60r2-32

OpenStudy (er.mohd.amir):

3r (r+8)-4(r+8) =3r2+24r-4r-32 =3r2+20r-32

OpenStudy (anonymous):

you multiply each of the first set with the second, like Er.Mohd.AMIR did.

OpenStudy (anonymous):

Oh...wow you guys made it simple to understand thank you so much!